OpenStudy (anonymous):

square root of 64x^6

jimthompson5910 (jim_thompson5910):

the square root of 64 is ______

OpenStudy (anonymous):

8

jimthompson5910 (jim_thompson5910):

the square root of x^6 is ______

OpenStudy (anonymous):

that i do not know.

jimthompson5910 (jim_thompson5910):

what you do is divide the exponent by 2

jimthompson5910 (jim_thompson5910):

in this case, the exponent is 6, so 6/2 = 3

jimthompson5910 (jim_thompson5910):

the square root of x^6 is x^3

jimthompson5910 (jim_thompson5910):

put this all together to get \[\large \sqrt{64x^6} = 8x^3\]

OpenStudy (anonymous):

so whatever the x^ is you / it by 2 for the square root

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

exactly

jimthompson5910 (jim_thompson5910):

another example \[\large \sqrt{x^{14}} = x^{14/2} = x^7\] \[\large \sqrt{x^{14}} = x^7\]

OpenStudy (anonymous):

even when the ^ is a square number like 16,36 ect. and what if its a odd number

OpenStudy (anonymous):

like ^25 would it be 12.5

jimthompson5910 (jim_thompson5910):

it would be \[\large \sqrt{x^{25}} = x^{25/2}\] or \[\large \sqrt{x^{25}} = x^{12.5}\] since 25/2 = 12.5

jimthompson5910 (jim_thompson5910):

alternatively, you can say this \[\large \sqrt{x^{25}} = x^{25/2}\] \[\large \sqrt{x^{25}} = x^{(24+1)/2}\] \[\large \sqrt{x^{25}} = x^{24/2}*x^{1/2}\] \[\large \sqrt{x^{25}} = x^{12}*x^{1/2}\] \[\large \sqrt{x^{25}} = x^{12}*\sqrt{x}\]

OpenStudy (anonymous):

ok what about (x+7)^2=-13

jimthompson5910 (jim_thompson5910):

what do you get when you square any number is that result ever negative?

OpenStudy (anonymous):

no so is that mean there is no answer

jimthompson5910 (jim_thompson5910):

there's no real numbered answer

jimthompson5910 (jim_thompson5910):

when you start to study imaginary and complex numbers, it will change how you answer but for now, there are no real solutions

OpenStudy (anonymous):

you wouldn't turn (x+7)^2=-13 into x^2+14x+49=-13

jimthompson5910 (jim_thompson5910):

you could, but that's going in the wrong direction

OpenStudy (anonymous):

if it was a positive number would you square that number and have x+7= the square root of 13

jimthompson5910 (jim_thompson5910):

yes this would work if 13 was positive, that would give you 2 real solutions

OpenStudy (anonymous):

then subtract the 7 from the square root of 13 to figure out x

jimthompson5910 (jim_thompson5910):

yeah if you had (x+7)^2 = 13 then you would get x = -7 + sqrt(13) or x = -7 - sqrt(13)

OpenStudy (anonymous):

so to the nearest thousandth would be 3.606 then i would -7+3.606 or -7+-3.606

jimthompson5910 (jim_thompson5910):

yeah but remember this only works if the 13 is positive in the original problem, the 13 is negative

OpenStudy (anonymous):

ok thanks for the help

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

if your still there how would i solve 7p^2=-9p-2 since the answer is a - dose that mean it dose not have a true number answer. I have to use the quadratic formula

jimthompson5910 (jim_thompson5910):

yeah i recommend using the quadratic formula

OpenStudy (anonymous):

if i remember it is x=-b square root of b^2-4ac and all of that over 2a

jimthompson5910 (jim_thompson5910):

-b +- sqrt(b^2 - 4ac) all over 2a

OpenStudy (anonymous):

once i get the sort(b^2-4ac) i then *the -b first then all of that is over 2a?

OpenStudy (anonymous):

right now i have -(-12)sqrt(144-4*3*-3 all over 6 the sqrt would be 48 so i have 12 sqrt 48 all over 6

jimthompson5910 (jim_thompson5910):

you are forgetting the plus or minus

jimthompson5910 (jim_thompson5910):

between -b and the square root

OpenStudy (anonymous):

12+or-the sqrt of 48 all over 2a

jimthompson5910 (jim_thompson5910):

you're solving 7p^2=-9p-2 still right? or is this a different problem?

OpenStudy (anonymous):

sorry my bad i am working on 3x^2-12x-3=0

jimthompson5910 (jim_thompson5910):

ok one sec

jimthompson5910 (jim_thompson5910):

this is what you should get \[\Large x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[\Large x = \frac{-(-12)\pm\sqrt{(-12)^2-4(3)(-3)}}{2(3)}\] \[\Large x = \frac{12\pm\sqrt{144-(-36)}}{6}\] \[\Large x = \frac{12\pm\sqrt{144+36}}{6}\] \[\Large x = \frac{12\pm\sqrt{180}}{6}\] \[\Large x = \frac{12+\sqrt{180}}{6} \ \text{or} \ x = \frac{12-\sqrt{180}}{6}\] \[\Large x = \frac{12+6\sqrt{5}}{6} \ \text{or} \ x = \frac{12-6\sqrt{5}}{6}\] \[\Large x = \frac{6(2+\sqrt{5})}{6} \ \text{or} \ x = \frac{6(2-\sqrt{5})}{6}\] \[\Large x = 2+\sqrt{5} \ \text{or} \ x = 2-\sqrt{5}\]

OpenStudy (anonymous):

how do I graph quadratic equation y=3x^2

jimthompson5910 (jim_thompson5910):

by plotting a series of points

jimthompson5910 (jim_thompson5910):

each point is found by plugging in values of x to get corresponding values of y

OpenStudy (anonymous):

all i need is 2 different values for x like 1 and 2?

OpenStudy (anonymous):

and figure what y= is x is 1 and if x is 2

jimthompson5910 (jim_thompson5910):

I would do more than 2 points

OpenStudy (anonymous):

would you do 4 points

jimthompson5910 (jim_thompson5910):

4 might be good

OpenStudy (anonymous):

when graphing y=(x-2)^2 I used the x value of 1,2,3,4 when i got the y value it was1,0,1,4 that made two parabolas one going facing up and one facing down? what dose this mean

jimthompson5910 (jim_thompson5910):

you should only get one parabola

jimthompson5910 (jim_thompson5910):

that is opening upward

OpenStudy (anonymous):

i see were i messed up, I need to plot them in order value